Consider the matrices A =   −2 7 1 3 4 1 8 1 5   ,B =   8 1 5 3 4 1 −2 7 1   , C =   −2 7 1 3 4 1 2 −7 3   . Find elementary matrices E1, E2 and E3 such that (5.1) E1A = B, (1) (5.2) E1B = A, (1) (5.3) E2A = C, (1) (5.4) E3C = A.


1
Expert's answer
2021-06-08T07:31:32-0400

(5.1)


"A=\begin{pmatrix}
 -2 & 7 & 1 \\
 3 & 4 & 1 \\
 8 & 1 & 5 \\
\end{pmatrix}, B=\begin{pmatrix}
 8 &1 & 5 \\
 3 & 4 & 1 \\
 -2 & 7 & 1 \\
\end{pmatrix}"

A matrix "A" switches all matrix elements on row "1" with their counterparts on row "3."

The corresponding elementary matrix is obtained by swapping row "1"  and row "3"  of the identity matrix.

"E_1=\begin{pmatrix}
 0 & 0 & 1 \\
 0 & 1 & 0 \\
 1 & 0 & 0 \\
\end{pmatrix}"

"E_1A=\begin{pmatrix}
 0 & 0 & 1 \\
 0 & 1 & 0 \\
 1 & 0 & 0 \\
\end{pmatrix}\begin{pmatrix}
 -2 & 7 & 1 \\
 3 & 4 & 1 \\
 8 & 1 & 5 \\
\end{pmatrix}"

"=\begin{pmatrix}
 8 &1 & 5 \\
 3 & 4 & 1 \\
 -2 & 7 & 1 \\
\end{pmatrix}=B"

"E_1=\begin{pmatrix}
 0 & 0 & 1 \\
 0 & 1 & 0 \\
 1 & 0 & 0 \\
\end{pmatrix}"



(5.2)

A matrix "B" switches all matrix elements on row "1" with their counterparts on row "3."

The corresponding elementary matrix is obtained by swapping row "1" and row "3" of the identity matrix.


"E_1=\begin{pmatrix}
 0 & 0 & 1 \\
 0 & 1 & 0 \\
 1 & 0 & 0 \\
\end{pmatrix}"

"E_1B=\begin{pmatrix}
 0 & 0 & 1 \\
 0 & 1 & 0 \\
 1 & 0 & 0 \\
\end{pmatrix}\begin{pmatrix}
 8 &1 & 5 \\
 3 & 4 & 1 \\
 -2 & 7 & 1 \\
\end{pmatrix}"

"=\begin{pmatrix}
 -2 & 7 & 1 \\
 3 & 4 & 1 \\
 8 & 1 & 5 \\
\end{pmatrix}=A"

"E_1=\begin{pmatrix}
 0 & 0 & 1 \\
 0 & 1 & 0 \\
 1 & 0 & 0 \\
\end{pmatrix}"


(5.3)


"A=\begin{pmatrix}
 -2 & 7 & 1 \\
 3 & 4 & 1 \\
 8 & 1 & 5 \\
\end{pmatrix}, C=\begin{pmatrix}
 -2 & 7 & 1 \\
 3 & 4 & 1 \\
 2 & -7 & 3 \\
\end{pmatrix}"

A matrix "C" is the matrix produced from "A" by adding "(-2)" times row "2" to row "3."

The corresponding elementary matrix is the identity matrix but with "(-2)" in the "(2,3)"

position.


"E_2=\begin{pmatrix}

 1 & 0 & 0 \\

 0 & 1 & 0 \\

 0 & -2 & 1\\

\end{pmatrix}"

"E_2A=\begin{pmatrix}
 1 & 0 & 0 \\
 0 & 1 & 0 \\
 0 & -2 & 1 \\
\end{pmatrix}\begin{pmatrix}
 -2 & 7 & 1 \\
 3 & 4 & 1 \\
 8 & 1 & 5 \\
\end{pmatrix}"

"=\begin{pmatrix}
 -2 & 7 & 1 \\
 3 & 4 & 1 \\
 2 & -7 & 3 \\
\end{pmatrix}=C"

"E_2=\begin{pmatrix}

 1 & 0 & 0 \\

 0 & 1 & 0 \\

 0 & -2 & 1\\

\end{pmatrix}"

(5.4)

A matrix "A" is the matrix produced from "C" by adding "(2)" times row "2" to row "3."

The corresponding elementary matrix is the identity matrix but with "(2)" in the "(2,3)"

position.


"E_3=\begin{pmatrix}

 1 & 0 & 0 \\

 0 & 1 & 0 \\

 0 & 2 & 1\\

\end{pmatrix}"

"E_3C=\begin{pmatrix}
 1 & 0 & 0 \\
 0 & 1 & 0 \\
 0 & 2 & 1 \\
\end{pmatrix}\begin{pmatrix}
 -2 & 7 & 1 \\
 3 & 4 & 1 \\
 2 & -7 & 3 \\
\end{pmatrix}"




"=\begin{pmatrix}
 -2 & 7 & 1 \\
 3 & 4 & 1 \\
 8 & 1 & 5 \\
\end{pmatrix}=A"

"E_3=\begin{pmatrix}

 1 & 0 & 0 \\

 0 & 1 & 0 \\

 0 & 2 & 1\\

\end{pmatrix}"